# %First␣names␣are␣abbreviated␣in␣the␣running␣head.%If␣there␣are␣more␣than␣two␣authors,␣’et␣al.’␣is␣used.http://www.docugami.com

### v1

@@ Abstract
Mathematical Reasoning Small Language Model Weak-Supervision.
@@ 1 Introduction
Recent developments have popularized Large Language Models (LLMs) [1, 24, 3] as potent tools for tackling complex mathematical reasoning problems [18]. Different strategies encompass an extensive range of techniques, including specialized models like [21, 31], Chain-of-Thought or Program-of-Thought [25, 4], or, tool-based compositional frameworks like [8, 14, 16, 27]. Progress in the study of mathematical reasoning with structured and unstructured data has been tangible thanks to various datasets and benchmarks, such as [5, 30, 17]. Tool-augmented LLMs are particularly prevalent in those benchmarks as they combine program assistance, data manipulation, natural language reasoning, and many other tools. These frameworks leverage the increasing planning capabilities of LLMs [26, 2] to orchestrate the tools, enabling multi-step problem-solving.
Figure 1: Example from FinQA with questions over tabular and textual data.
Some studies focus on prompt engineering methods to get the best from closed-source models like GPT-4 [22, 16, 25, 4]. Others involve fine-tuning multiple specialized Small Language Models (SLMs) in a teacher-student fashion with data annotated automatically by larger models [14, 8], or extensively by human experts [31, 30, 5]. Other methods focus on self-improvement, with either training verifiers [10], generators for data augmentation [28], or multi-sampling techniques [29]. The diversity in the types of approaches underscores the potential of language models in enhancing reasoning capabilities.
However, each approach faces dependencies or limitations, summarized in Table 1. Prompt engineering is labor-intensive, usually requires hand-crafting efforts, and may not fully capture data complexity and diversity. Closed-source models raise privacy concerns, especially with business documents, and present scalability issues due to their cost and latency, whether for querying models directly or generating high-quality reasoning data. Moreover, using teacher models to generate reasoning steps is often data-inefficient [10], discarding incorrect solutions, and human expert annotation [31] is a costly alternative. To overcome this issue, various self-improvement methods have been developed. However, most of them require multiple samplings and/or additional models [13, 10], leading to an increase in computational and time costs.
Inspired by those challenges, this work introduces a novel approach for mathematical reasoning over tabular data. The proposed method aims to reduce both prompt engineering efforts, reliance on larger or external models, privacy concerns, and computational inefficiencies while maintaining comparable performances to existing methods.
Table 1: Overview of LLM reasoning frameworks over tabular data. PE: Prompt Engineering; FT: Fine-Tuning.
Method Tool-use GPT-4 reliance Human-experts Datasets used
@@
The main contributions of this work are: 1) a tool-augmented framework that composes shared tools from diverse datasets; 2) a weak-supervised fine-tuning for the tools from self-generated correct reasoning trajectories; 3) a weak-supervised preference optimization to align the tools using all updated reasoning trajectories.
@@ 2.1 Datasets for Mathematical Reasoning
Evaluating LLMs’ mathematical reasoning capabilities is usually focused on datasets like MATH [9] and GSM8K [6], which explore mathematical problem-solving. While these benchmarks span a broad domain of mathematical subjects, the data format is limited to textual word problems. In contrast, other benchmarks like FinQA [5] and TAT-QA [30] represent more complex and real-world reasoning scenarios. These datasets include tabular and textual data, requiring reasoning over financial reports with intricated contexts. Moreover, TabMWP [17] expands to mathematical word problems over different subjects. Those three datasets need multi-step reasoning, information extraction, data manipulation, tabular and contextual understanding and numerical calculations.
@@ 2.2 Tool-Augmented Frameworks
Multi-step reasoning has emerged as an approach for mathematical reasoning, particularly over tabular data. Diverse prompting techniques [25, 4, 15, 22] have been explored to encourage stepwise and critical thinking, traditionally relying on closed-source LLMs. Other frameworks adopt LLMs as tools with a dedicated planning: both TORA [8] and Husky [14] fine-tune open-source LLMs with training data generated by GPT-4, while Chameleon [16] relies on prompt engineering, directly backed up by ChatGPT/GPT-4. TAT-LLM [31] relies on a fixed stepwise pipeline for reasoning decomposition, with a fine-tuned model trained on data heavily annotated by human experts. Those fine-tuning approaches rely on closed-source models or human-annotations to generate high-quality synthetic data to train on.
@@ 2.3 Alignment Methods
To align LLMs on downstream tasks, methods have progressed from Reinforcement Learning from Human Feedback (RLHF) [19] to Direct Preference Optimization (DPO) [20], and more recently Kahneman-Tversky Optimization (KTO) [7]. DPO offers a cost-effective alternative to RLHF by eliminating the need for a separately trained reward model. Instead DPO relies on a preference dataset with positive and negative completions for given inputs. In contrast to traditional methods that optimize the log-likelihood of preferences, KTO focuses on maximizing generation utility. KTO is also less data-intensive than DPO as it leverages binary preference data, i.e., completions paired with boolean labels indicating the correctness of the completion. Therefore, it is particularly well-suited for downstream tasks following Supervised Fine-Tuning (SFT).
@@ 3.1 Motivations
MATATA is a novel cost-effective method for training LLM agents designed to tackle tabular data problems through reasoning, planning, and tool use. This work explores the following questions: Can high performance be achieved while reducing reliance on external models, data, and extensive prompt engineering? Can a tool-augmented SLM framework for mathematical reasoning self-improve through weak supervision?
@@ 3.2.1 MATATA Compositional Reasoning
3.2.2 MATATA Weak-Supervision Stages
@@
Our work redefines the teacher-student paradigm as the same model generates the trajectories used for weak supervision fine-tuning. The first training stage is SFT [19], where each tool is fine-tuned on prompt/completion pairs from valid trajectories generated by the baseline system. After the SFT stage, tools are aligned with KTO [7]. Using KTO is motivated by weak supervision, as it only requires a binary signal (correct/incorrect). KTO does not require multiple samplings for a given problem, unlike DPO [20], or additional models like verifiers or generators [10], or bootstrapping methods [29]. At tool level, for each prompt/completion pair, the preference label is the weak signal from the final answer, without requiring step-by-step annotations. Using all invalid trajectories during the alignment phase could help increase domain knowledge in areas uncovered with SFT, countering the tendency of other methods to improve mostly in the framework’s areas of proficiency.
@@ 4 Experiments
4.1 Training Details
Experiments are conducted on FinQA [5], TAT-QA [30], and TabMWP [17]. The selected tools for this experiment are inspired by Chameleon [16]. The following tools are shared across datasets: Program Generator, Solution Generator (Natural Language reasoner), Context Extractor, Row/Column Extractor, and Knowledge Retrieval. The task-specific tools are: a unique Planner for each dataset, and a Span Extractor for TAT-QA only. This modular design enables flexible handling of diverse datasets, allowing to add and share new tools when incorporating additional datasets.
Table 2: Accuracy/Exact-Match scores across evaluation datasets. PE: Prompt Engineering; FT: Fine-Tuning. Best overall scores underlined, best section scores bolded.
Framework Model Method FinQA TAT-QA TabMWP
Acc. EM Acc.
@@ 4.1 Training Details
Chameleon [16] ChatGPT PE - - 93.28
GPT-4 PE - - 98.78
EEDP [22] ChatGPT PE 61.88 79.73 -
GPT-4 PE 76.05 88.67 -
PoT [4] PoT-SC-Codex PE 68.1 70.2 81.8
@@
FinQANet [5] RoBERTa FT 61.24 - -
TagOP [30] RoBERTa FT - 50.1 -
TORA [8] Code-LLama-34B FT - - 70.5
Llama2-70B FT - - 74.0
@@
TAT-LLM [31] Llama2-7B FT 65.13 76.4 -
Llama2-13B FT 71.93 77.51 -
Llama2-70B FT 76.81 81.42 -
@@
Train sets are used to generate reasoning trajectories and train the tools. The best model is selected with respect to scores on validation sets. Results on the test sets are reported in Table  2 using the datasets metrics: Accuracy or Exact-Match. All datasets and splits are used entirely, as in their respective papers.
Two light-weight language models are used as baseline and for fine-tuning: microsoft/Phi-3-mini-4k-instruct 3.8B [1] and mistralai/Ministral-8B-Instruct-2410 [23]. All trainings are conducted on a single NVIDIA A100 GPU, for a combined total of about 128 GPU hours. Throughout the training process, one adapter is trained for each tool $t$ in the framework $M_{t}\rightarrow M^{SFT}_{t}\rightarrow M^{SFT+KTO}_{t}$. Both SFT and KTO adapter fine-tuning use the following hyper-parameters: a learning rate of 1e-5 and a batch size of 32 over 10 epochs, with a LoRA $r$ of 64, $\alpha$ of 32 and dropout of 0.05. KTO uses the default $\beta$ parameters at 0.1, and task-specific weights are applied to address class imbalance between valid and invalid trajectories.
@@ 4.2 Experimental Results
Table 3: Evolution of MATATA-3.8B and MATATA-8B Accuracy and EM over training stages; with one vs. all training set(s).
Method Training dataset(s) FinQA TAT-QA TabMWP
Phi3-mini 3.8B baseline - 47.06 43.05 83.87
@@
The experimental results in Table 2 show the performances of MATATA-Phi3-mini 3.8B and MATATA-Ministral 8B, respectively referred to as MATATA-3.8B and MATATA-8B. MATATA outperforms Husky, which relies on training data annotated by GPT-4, and PoT-Codex, a GPT-3 variant, on all 3 datasets. Both RoBERTa-based baseline models for FinQA and TAT-QA are also surpassed. On TAT-QA, MATATA-8B ranges between TAT-LLM-13B and TAT-LLM-70B, while TAT-LLM uses human-expert annotations for training. The gap between MATATA-8B and TAT-LLM-70B is of only 3.61%, despite the model size decrease. On FinQA, MATATA-8B is above all methods, and MATATA-3.8B ranges between TAT-LLM-7B and -13B, and between ChatGPT and GPT-4 with EEDP-prompting, while being much smaller than those models. On TabMWP, MATATA-8B exceeds all methods not using GPT-4, ranking 2nd on the TabMWP leaderboard, with only a 0.66% difference behind Chameleon GPT-4. For its range of parameters and size, MATATA competes with models fine-tuned on higher-quality reasoning data, larger, or extensively prompted. This also confirms the effectiveness of our method, intuiting that fine-tuning a larger model with weak-supervision could further improve those scores.
To investigate knowledge transfer across datasets with shared tools, two training strategies are explored: (1) training the tools on separate datasets, and (2) training shared tools on the combination of datasets. Results reported in Table 3 show that training tools with weak-trajectories from all datasets outperforms single dataset training, in both SFT and KTO stages. This improvement can be attributed to a better coverage over trajectories of each tool, with the increase and diversification of training data. It also confirms the reusability and scalability of MATATA: the same framework could be enhanced with other datasets and more shared tools, bringing potential improvements to existing results. Also, the performance increase from baseline models to fine-tuned ones is between 12.79% and 31.39% for the 3.8B model, and between 17.23% and 31.39% for the 8B model. Although the baseline model receives minimal initial guidance through low-effort prompt engineering, applying weak-supervision with SFT and KTO enables tools to naturally adopt their expected behaviors, improving the baseline performances significantly.
@@ 5 Conclusion
This work explore a novel method for language agents to tackle tabular data problems using only SLMs, without relying neither on larger and/or close-source models, or external data. This is essential for sensitive business documents, such as those in FinQA or TAT-QA, as submitting those to closed-source LLMs can raise security concerns. With minimal efforts spent on hand-crafting the prompts, human labor is also reduced, adapting the framework to its users rather than the opposite. Competitive results validate that tool-augmented SLMs with weak-supervision can meet performances of existing reasoning methods. This work serves as a step for future developments in creating small, reliable, and secure locally-operated SLM frameworks capable of complex reasoning tasks.
@@
5.0.1 Acknowledgements
This work was partially supported by Mathematics of Information Technology and Complex Systems (MITACS) Accelerate, grant IT41737.

### v2

@@ Abstract
Mathematical Reasoning Small Language Model Weak Supervision.
@@ 1 Introduction
Recent developments have popularized Large Language Models (LLMs) [1, 24, 3] as potent tools for tackling complex mathematical reasoning problems [18]. Different strategies encompass an extensive range of techniques, including specialized models like [21, 31], Chain-of-Thought or Program-of-Thought [25, 4], or, tool-based compositional frameworks like [8, 14, 16, 27]. Progress in the study of mathematical reasoning with structured and unstructured data has been measured thanks to various datasets and benchmarks, such as [5, 30, 17]. Tool-augmented LLMs are particularly prevalent in those benchmarks as they combine program assistance, data manipulation, natural language reasoning, and many other tools. These frameworks leverage the increasing planning capabilities of LLMs [26, 2] to orchestrate tools, enabling multi-step problem solving.
Figure 1: Example from FinQA [5] with questions over tabular and textual data.
Recent studies focus on prompt engineering methods to get the best from closed-source models like GPT-4 [22, 16, 25, 4]. Others involve fine-tuning multiple specialized Small Language Models (SLMs) in a teacher-student fashion with data annotated automatically by larger models [14, 8], or extensively by human experts [31, 30, 5]. Further methods focus on self-improvement, with either training verifiers [10], generators for data augmentation [28], or multi-sampling techniques [29]. The diversity in the types of approaches underscores the potential of language models in enhancing reasoning capabilities.
However, each approach faces dependencies or limitations, summarized in Table 1. Prompt engineering is labor-intensive, usually requires hand-crafting efforts, and may not fully capture data complexity and diversity. Closed-source models raise privacy concerns, especially with business documents, and present scalability issues due to their cost and latency, whether for querying models directly or generating high-quality reasoning data. Moreover, using teacher models to generate reasoning steps is often data-inefficient [10], discarding incorrect solutions, and human expert annotation [31] is a costly alternative. To overcome this issue, various self-improvement methods have been developed. Although, most of them require multiple samplings and/or additional models [13, 10], leading to an increase in computational and time costs.
Motivated by those challenges, this work introduces a novel approach for mathematical reasoning over tabular data. The proposed method aims to reduce both prompt engineering efforts, reliance on larger or external models, privacy concerns, and computational inefficiencies while maintaining comparable performances to existing methods.
Table 1: Overview of LLM reasoning frameworks over tabular data, with their learning method, Prompt Engineering (PE) or Fine-Tuning (FT), if they are tool-augmented, their reliance on GPT-4, their use of human experts and their evaluation datasets.
Tool- Uses Human
Framework Method aug. GPT-4 experts Datasets used
@@
The main contributions of this work are: 1) a tool-augmented framework that composes shared tools from diverse datasets; 2) a weak-supervised fine-tuning for the tools from self-generated correct reasoning trajectories; 3) a weak-supervised preference optimization to align the tools using all available reasoning trajectories.
@@ 2.1 Datasets for Mathematical Reasoning
Evaluating LLMs’ mathematical reasoning capabilities is usually focused on datasets like MATH [9] and GSM8K [6], which explore mathematical problem-solving. While these benchmarks span a broad domain of mathematical subjects, the data format is limited to textual word problems. In contrast, other benchmarks like FinQA [5] and TAT-QA [30] represent more complex and real-world reasoning scenarios. These datasets include tabular and textual data, requiring reasoning over financial reports with intricate contexts. Moreover, TabMWP [17] expands to mathematical word problems over different subjects. Those three datasets need multi-step reasoning, information extraction, data manipulation, tabular and contextual understanding and numerical calculations.
@@ 2.2 Tool-Augmented Frameworks
Multi-step reasoning has emerged as an approach for mathematical reasoning, particularly over tabular data. Diverse prompting techniques [25, 4, 15, 22] have been explored to encourage stepwise and critical thinking, traditionally relying on closed-source LLMs. Other frameworks adopt LLMs as tools with a dedicated planning: both TORA [8] and Husky [14] fine-tune open-source LLMs with training data generated by GPT-4, while Chameleon [16] relies on prompt engineering, directly backed up by ChatGPT/GPT-4. TAT-LLM [31] relies on a fixed stepwise pipeline for reasoning decomposition, with a fine-tuned model trained on data annotated by human experts. Those fine-tuning approaches rely on closed-source models or human-annotations to generate high-quality synthetic data to train on.
@@ 2.3 Alignment Methods
To align LLMs on downstream tasks, methods have progressed from Reinforcement Learning from Human Feedback (RLHF) [19] to Direct Preference Optimization (DPO) [20], and more recently Kahneman-Tversky Optimization (KTO) [7]. DPO offers a cost-effective alternative to RLHF by eliminating the need for a separately trained reward model. Instead, DPO relies on a preference dataset with positive and negative completions for given inputs. KTO focuses on maximizing generation utility, in contrast to traditional methods that optimize the log-likelihood of preferences. KTO is also less data-intensive than DPO as it leverages binary preference data, i.e., completions paired with Boolean labels indicating the correctness of the completion. Therefore, it is particularly well-suited for downstream tasks following Supervised Fine-Tuning (SFT).
@@ 3.1 Motivations
MATATA is a novel cost-effective method for training LLM agents designed to tackle tabular data problems through reasoning, planning, and tool use. This work explores the following questions:
  • Can high performance be achieved while reducing reliance on external models, data, and extensive prompt engineering?
  • Can a tool-augmented SLM framework for mathematical reasoning self-improve through weak supervision?
@@ 3.2.1 MATATA Compositional Reasoning
3.2.2 MATATA Weak Supervision Stages
@@
Our work redefines the teacher-student paradigm as the same model generates the trajectories used for weak supervision fine-tuning. The first training stage is SFT [19], where each tool is fine-tuned on prompt/completion pairs from valid trajectories generated by the baseline system.
After the SFT stage, tools are aligned with KTO [7]. Using KTO is motivated by weak supervision, as it only requires a binary signal (correct/incorrect). KTO does not require multiple samplings for a given problem, unlike DPO [20], or additional models like verifiers or generators [10], or bootstrapping methods [29]. At tool level, for each prompt/completion pair, the preference label is the weak signal from the final answer, without requiring step-by-step annotations. Using all invalid trajectories during the alignment phase could help increase domain knowledge in areas uncovered with SFT, countering the tendency of other methods to improve mostly in the framework’s areas of proficiency.
@@ 4 Experiments
Table 2: Accuracy (acc.) or Exact-Match (EM) scores across evaluation datasets. Prompt Engineering (PE); Fine-Tuning (FT). Best overall scores appear underlined, best section scores appear in bold.
FinQA TAT-QA TabMWP
Framework Model Method Acc. EM Acc.
@@
Chameleon [16] ChatGPT PE 93.28
GPT-4 PE 98.78
EEDP [22] ChatGPT PE 61.88 79.73
GPT-4 PE 76.05 88.67
PoT [4] SC-Codex PE 68.1 70.2 81.8
GPT-4 PE 74.0
@@
FinQANet [5] RoBERTa FT 61.24
TagOP [30] RoBERTa FT 50.1
TORA [8] Code-LLama-34B FT 70.5
Llama2-70B FT 74.0
@@
TAT-LLM [31] Llama2-7B FT 65.13 76.4
Llama2-13B FT 71.93 77.51
Llama2-70B FT 76.81 81.42
@@
4.1 Training Details
Experiments are conducted on the FinQA [5], TAT-QA [30], and TabMWP [17] datasets. The selected tools for this experiment are inspired by Chameleon [16]. The following tools are shared across datasets: Program Generator, Solution Generator (Natural Language reasoner), Context Extractor, Row/Column Extractor, and Knowledge Retrieval. The task-specific tools are: a unique Planner for each dataset, and a Span Extractor for TAT-QA only. This modular design enables flexible handling of diverse datasets, allowing to add and share new tools when incorporating additional datasets.
Train sets are used to generate reasoning trajectories and train the tools. The best model is selected with respect to scores on the validation sets. The results on the test sets are reported in Table 2 using the relative metric for each dataset, either Accuracy or Exact Match. All datasets and splits are used entirely, as in their respective papers.
Two light-weight language models are used as baselines and for fine-tuning: microsoft/Phi-3-mini-4k-instruct 3.8B [1] and mistralai/Ministral-8B-Instruct-2410 [23]. All trainings are conducted on a single NVIDIA A100 GPU, for a combined total of about 128 GPU hours. Throughout the training process, one adapter is trained for each tool $t$ following: $M_{t}\rightarrow M^{SFT}_{t}\rightarrow M^{SFT+KTO}_{t}$. Both SFT and KTO adapter fine-tuning use the following hyperparameters: a learning rate of 1e-5 and a batch size of 32 over 10 epochs, with a LoRA $r$ of 64, $\alpha$ of 32 and a dropout of 0.05. KTO uses the default $\beta$ parameters at 0.1, and task-specific weights are applied to address class imbalance between valid and invalid trajectories.
@@ 4.2 Experimental Results
Table 3: Evolution of MATATA-3.8B and MATATA-8B accuracy or exact match over fine-tuning stages; with one vs. all training set(s).
Fine-tuned FinQA TAT-QA TabMWP
Method on Acc. EM Acc.
Phi3-mini 3.8B baseline N/A 47.06 43.05 83.87
@@
The experimental results in Table 2 show the performance of MATATA-Phi3-mini 3.8B and MATATA-Ministral 8B, respectively referred to as MATATA-3.8B and MATATA-8B. Results show that MATATA outperforms Husky, which relies on training data annotated by GPT-4, and PoT-Codex, a GPT-3 variant, on all 3 datasets. Both RoBERTa-based baseline models for FinQA and TAT-QA are also surpassed by MATATA.
On TAT-QA, MATATA-8B ranges between TAT-LLM-13B and TAT-LLM-70B, while TAT-LLM uses human-expert annotations for training. Notably, the gap between MATATA-8B and TAT-LLM-70B is only 3.61%, despite the decrease in model size. On FinQA, MATATA-8B is above all methods, and MATATA-3.8B ranges between TAT-LLM-7B and 13B, and between ChatGPT and GPT-4 with EEDP-prompting, while being much smaller than those models. On TabMWP, MATATA-8B exceeds all methods not using GPT-4, ranking 2nd on the TabMWP leaderboard, with only a 0.66% difference behind Chameleon which relies on the massively larger GPT-4. For its range of parameters and size, MATATA competes with models fine-tuned on higher-quality reasoning data, larger, or extensively prompted. This also confirms MATATA’s effectiveness, intuiting that fine-tuning a larger model with weak supervision could further improve those scores.
To investigate knowledge transfer across datasets with shared tools, two training strategies are explored: (1) training the tools on separate datasets, and (2) training shared tools on the combination of datasets. Results reported in Table 3 show that training tools with weak-trajectories from all datasets outperforms single dataset training, in both SFT and KTO stages. This improvement can be attributed to a better coverage over trajectories of each tool, with the increase and diversification of training data. It also confirms the reusability and scalability of MATATA: the same framework could be enhanced with other datasets and more shared tools, bringing potential improvements to existing results. Similarly, the performance increase from baseline models to fine-tuned ones is between 12.79% and 31.39% for the 3.8B model, and between 17.23% and 31.39% for the 8B model. Although the baseline model receives minimal initial guidance through low-effort prompt engineering, applying weak supervision with SFT and KTO enables tools to naturally adopt their expected behaviors, significantly improving the baseline performance.
@@ 5 Conclusion
This work explores a novel method for language agents to tackle tabular data problems using only SLMs, without relying neither on larger and/or close-source models, or external data. This is essential for sensitive business documents, exemplified by the ones present in FinQA or TAT-QA, as submitting those to closed-source LLMs -in particular when they are hosted as a cloud-based infrastructure- can raise security and privacy concerns. Another improvement is derived from minimal effort dedicated to hand-crafting the prompts. This has reduced the human labor, as the framework adapts to its users, rather than the opposite. Competitive results validate that tool-augmented SLMs with weak supervision can meet performances of existing reasoning methods. This work serves as a step towards future developments in creating small, reliable, and secure locally-operated SLM frameworks capable of complex reasoning tasks.
@@
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Additional Experiment Details
Additional details of the experiments over training data are summarized in Table 4. Correct examples from baselines are used for SFT fine-tuning, while correct and incorrect examples from SFT fine-tuning are used for KTO fine-tuning.
Table 4: Performance on training data of MATATA-3.8B and MATATA-8B over different configurations of the training sets.
Training Examples
Dataset Model set Total Correct Incorrect
FinQA Phi3-mini 3.8B baseline N/A 6251 2992 3259
+ weak-sup SFT One 6251 3756 2495
All 6251 4148 2103
TabMWP Phi3-mini 3.8B baseline N/A 23059 19808 3251
+ weak-sup SFT One 23059 21318 1741
All 23059 22280 779
TAT-QA Phi3-mini 3.8B baseline N/A 13205 6283 6922
+ weak-sup SFT One 13205 9122 4083
All 13205 9970 3235
FinQA Ministral 8B baseline N/A 6251 3434 2817
+ weak-sup SFT One 6251 4828 1423
All 6251 4900 1351
TabMWP Ministral 8B baseline N/A 23059 19019 4040
+ weak-sup SFT One 23059 20887 2172
All 23059 22436 623
TAT-QA Ministral 8B baseline N/A 13205 6705 6500
+ weak-sup SFT One 13205 10641 2564
All 13205 11045 2160